# Hannes Alfvéns väg 12, SE-106 91 Stockholm, Sweden

### v1

@@ Abstract
We present new black hole solutions to the low-energy effective action of string theory. We introduce three- and four-dimensional solutions that are rotating, asymptotically flat, and exhibit a linear dilaton vacuum. These solutions cannot be overspun, i.e., do not have an extremality condition, akin to higher-dimensional Myers-Perry black holes with one rotational angle. Studying their thermodynamics reveals that the temperature associated to these solutions does not depend on the black hole mass, similar to the Witten black hole. We also find that their asymptotic symmetry group is more stringent than the BMS group. We consider the charged generalisations for these black holes, which introduces closed timelike curves within the inner horizon. We show that these black holes can be derived from the large-$d$ limit of the Myers-Perry black hole. As such we advocate that large-$d$ can provide a useful vantage point to interpret the here introduced black holes, as well as more generally a way to generate new effective field theories and corresponding non-trivial solutions.
@@ 1 Introduction
Linear dilaton spacetimes notably feature in non-AdS/CFT realisations of holography where in particular the $D=2$ version of the low-energy effective action has been explored, see e.g. Aharony et al. (1998); Maldacena and Strominger (1997). We also present an uplift of this three-dimensional solution to four dimensions, which requires and extra $\ell^{2}(1+\alpha^{2}\cos^{2}\theta_{0})\text{d}\theta^{2}$ term in the metric and yields a cylindrical black hole solution with the same qualitative features. We can extend these solutions to an electrically charged black hole by adding a non-minimally coupled Maxwell field to the action.
@@
We present our work in the following order. We start by explicitly showing how the here presented $D=3$ and $D=4$ solutions solve the equations of motion of the low-energy effective action in Section 2. In Section 3 we study properties of these solutions, such as their thermodynamics. We elaborate on the large-$d$ pedigree in Section 4. A discussion of the results and ideas for future work are presented in Section 5.
@@ 3.2 Charges and thermodynamics
where $\star$ is the Hodge dual operator, $\sigma^{(t)}=g_{tt}\text{d}t+g_{t\phi}\text{d}\phi$ is the dual form of the timelike Killing vector $(\partial_{t})^{\mu}$, and $\sigma^{(t)}=g_{t\phi}\text{d}t+g_{\phi\phi}\text{d}\phi$ is the dual form of the spacelike Killing vector $(\partial_{\phi})^{\mu}$. Notice that the normalisation constants are different in the mass and angular momentum formulas, which is done to match the Komar mass and angular momentum with the mass and angular momentum in the non-relativistic (Newtonian) limit. Substituting with the three-dimensional charged metric (26), we get the Komar mass
@@ 4.2 Large-$d$ geometry of Myers-Perry
In order to obtain a non-trivial near-horizon description in the large-$d$ limit, we introduce the parametrisation $\mu=r_{0}^{d-1}$, where $r_{0}$ is a length scale. From $\Delta(r_{h})=0$, see (74), we establish
@@
$\frac{\mu}{r^{d-3}\Sigma}\stackrel{{\scriptstyle d\to\infty}}{{=}}$ $\penalty 10000\ e^{-w}\frac{1+\alpha^{2}}{1+\alpha^{2}\cos^{2}\theta_{0}}\,$,
@@ 5 Discussion
In this article we presented new rotating black hole solutions to the low-energy effective action of string theory, which can be viewed as a model of dilaton gravity. We focussed on a three-dimensional model as well as its four-dimensional extension. We find that its thermodynamics exhibit a mixture of higher dimensional Myers-Perry qualities such as having no extremal bound associated to rotation and two-dimensional Witten black hole features such as a temperature that does not depend on the black hole mass. The tool we used to track down the above solutions was studying the large-$d$ behaviour of Myers-Perry black holes. We furthermore also showed the extensions of the solutions for when we add an electric charge. Finally, we point out that all of the here presented models allow for a linear dilaton vacuum and the asymptotic symmetries of these models turn out to be more stringent than their Einstein gravity equivalents.
As an outlook to future work, we recall that the Witten black hole, which the here studied solutions closely mirror, has for example been important for studying dualities in string theory Fateev et al. (2000); Alvarez et al. (1994), studying connections to little string theory Aharony et al. (1998); Maldacena and Strominger (1997), evaporation Callan et al. (1992); Russo et al. (1992), the Page curve Gautason et al. (2020), and holographic complexity Schneiderbauer and others (2020); Schneiderbauer et al. (2020). It would be interesting to further explore our solutions along those lines. More along the direction of studying Myers-Perry black holes, it can be insightful to use our results to, e.g., study black holes with different topologies and instabilities such as the Gregory-Laflamme type Gregory and Laflamme (1993).
@@ Acknowledgements.
We thank Vyshnav Mohan, Bo Sundborg, and Lárus Thorlacius for stimulating discussions in the context of this article. We want to especially thank Roberto Emparan for suggesting the current narrative of this work and insightful comments. The work of WS is supported by a Starting Grant 2023-03373 from the Swedish Research Council and the Olle Engkvists Stiftelse. PT is supported by a Nordita visiting PhD student fellowship, Primus grant PRIMUS/23/SCI/005 and GAUK 425425 from Charles University.
@@ Appendix A Alternative derivation for rotation and temperature
Let us now compute the temperature and angular velocity. We assume the following near-horizon form:

### v2

@@ Abstract
We present new black hole solutions to the low-energy effective action of string theory. We introduce three- and four-dimensional solutions that are rotating, asymptotically flat, and exhibit a linear dilaton vacuum. We also introduce higher-dimensional generalisations of these black holes that possess multi angular momentum-like charges. These solutions cannot be overspun, i.e., do not have an extremality condition, akin to some higher-dimensional Myers-Perry black holes. Studying their thermodynamics reveals that the temperature associated to these solutions does not depend on the black hole mass, similar to the Witten black hole. We also find that their asymptotic symmetry group is more stringent than the BMS group. We consider the charged generalisations for these black holes, which introduces closed timelike curves within the inner horizon. We show that these black holes can be derived from the large-$d$ limit of the Myers-Perry black hole. As such we advocate that large-$d$ can provide a useful vantage point to interpret the here introduced black holes, as well as more generally a way to generate new effective field theories and corresponding non-trivial solutions.
@@ 1 Introduction
Linear dilaton spacetimes notably feature in non-AdS/CFT realisations of holography where in particular the $D=2$ version of the low-energy effective action has been explored, see e.g. Aharony et al. (1998); Maldacena and Strominger (1997). We also present an uplift of this three-dimensional solution to four dimensions, which requires and extra $\ell^{2}(1+\alpha^{2}\cos^{2}\theta_{0})\text{d}\theta^{2}$ term in the metric and yields a cylindrical black hole solution with the same qualitative features. We can extend these solutions to an electrically charged black hole by adding a non-minimally coupled Maxwell field to the action. Finally, we also present a solution that possesses multi angular momentum-like charges and is valid for any $D>2$.
@@
We present our work in the following order. We start by explicitly showing how the here presented $D=3$ and $D=4$ solutions solve the equations of motion of the low-energy effective action in Section 2. In Section 3 we study properties of these solutions, such as their thermodynamics. We elaborate on the large-$d$ pedigree in Section 4. In Section 5 we present a black hole solution that possesses multi angular momentum-like charges and discuss some of its properties. A discussion of the results and ideas for future work are presented in Section 6.
@@ 3.2 Charges and thermodynamics
where $\star$ is the Hodge dual operator, $\sigma^{(t)}=g_{tt}\text{d}t+g_{t\phi}\text{d}\phi$ is the dual form of the timelike Killing vector $(\partial_{t})^{\mu}$, and $\sigma^{(\phi)}=g_{t\phi}\text{d}t+g_{\phi\phi}\text{d}\phi$ is the dual form of the spacelike Killing vector $(\partial_{\phi})^{\mu}$. Notice that the normalisation constants are different in the mass and angular momentum formulas, which is done to match the Komar mass and angular momentum with the mass and angular momentum in the non-relativistic (Newtonian) limit. Substituting with the three-dimensional charged metric (26), we get the Komar mass
@@ 4.2 Large-$d$ geometry of Myers-Perry
In order to obtain a non-trivial near-horizon description in the large-$d$ limit, we introduce the parametrisation $\mu=r_{0}^{d-3}$, where $r_{0}$ is a length scale. From $\Delta(r_{h})=0$, see (74), we establish
@@
$\frac{\mu}{r^{d-5}\Sigma}\stackrel{{\scriptstyle d\to\infty}}{{=}}$ $\penalty 10000\ e^{-w}\frac{1+\alpha^{2}}{1+\alpha^{2}\cos^{2}\theta_{0}}\,$,
@@ 4.3 Reduction at the level of the action
5 A black hole possessing multi angular momentum-like charges
Generalising the large-$d$ tools constructed in the previous paper we can derive a solution to the general $D$-dimensional low-energy effective action (2) that possesses multi angular momentum-like charges. The Einstein’s equation and dilaton equation of motion associated to (2) are:
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+2g_{\mu\nu}(\partial\Phi)^{2}+2\nabla_{\mu}\nabla_{\nu}\Phi-2g_{\mu\nu}\Box\Phi-2\lambda^{2}g_{\mu\nu}=0\,$, (103)
$R-4(\partial\Phi)^{2}+4\lambda^{2}+4\Box\Phi=0\,.$ (104)
We claim that for any dimension $D\geq 2$, the equations of motion written above are solved by the following geometry [^4] A Mathematica notebook with more details is available upon request.
$\text{d}s^{2}_{D}=$ $\penalty 10000\ -\text{d}t^{2}+\mathcal{A}_{D}(1-\mathcal{F})\left(\text{d}t+\sum_{\mathclap{i=1}}^{D-2}\ell\alpha_{i}\mu_{0}{}_{i}^{2}\text{d}\phi_{i}\right)^{2}+\frac{\ell^{2}\text{d}w^{2}}{\mathcal{A}_{D}\mathcal{F}}$ (105)
$\penalty 10000\ +\ell^{2}\sum_{\mathclap{i=1}}^{D-2}(1+\alpha_{i}^{2})\mu_{0}{}_{i}^{2}\text{d}\phi_{i}^{2}\,$,
with
$\mathcal{F}=1-me^{-w}\,,\quad 4\lambda^{2}=\frac{\mathcal{A}_{D}}{\ell^{2}}\,,\quad\mathcal{A}_{D}=\left(1-\sum^{D-2}_{i=1}\frac{\alpha_{i}^{2}\mu_{0}{}_{i}^{2}}{1+\alpha^{2}_{i}}\right)^{-1}\,$, (106)
and supplemented with a dilaton solution that also exhibits an asymptotically linear dilaton vacuum
$\Phi=-\frac{1}{2}w\,.$ (107)
In the above solution $\alpha_{i}$ are the parameters proportional to the angular momenta associated to the angles $\phi_{i}$. The parameters $\mu_{0}{}_{i}>0$ encode the relative warping of the asymptotic spacetime. The Ricci scalar is given by
$R=\frac{\mathcal{A}_{D}}{\ell^{2}}e^{-w}\,$, (108)
which reveals a curvature singularity for $w\to$-$\infty$ and an asymptotic flat space behaviour.
Due to a similar emblackening factor as the single-angle rotating case the thermodynamic behaviour is qualitatively the same, notably exhibiting no Kerr-like bound. Its asymptotic symmetries are
$\text{BMS}_{2}\times\underbrace{U(1)\times\cdots\times U(1)}_{D-2\ \text{times}}\,$, (109)
which is more stringent than the full BMSD group. Finally, contrasting the Myers-Perry solution that has $\lfloor(D-1)/2\rfloor$ angular momenta, the here treated solution allows for D2D-2 angular momenta-like charges.
@@ 6 Discussion
In this article we presented new rotating black hole solutions to the low-energy effective action of string theory, which can be viewed as a model of dilaton gravity. We focussed on a three-dimensional model as well as its four-dimensional extension. We find that its thermodynamics exhibit a mixture of higher dimensional Myers-Perry qualities such as having no extremal bound associated to rotation and two-dimensional Witten black hole features such as a temperature that does not depend on the black hole mass. The tool we used to track down the above solutions was studying the large-$d$ behaviour of Myers-Perry black holes. We furthermore also showed the extensions of the solutions for when we add an electric charge or extra angular momentum-like charges for arbitrary dimensions.Finally, we point out that all of the here presented models allow for a linear dilaton vacuum and the asymptotic symmetries of these models turn out to be more stringent than their Einstein gravity equivalents.
As an outlook to future work, we recall that the Witten black hole, which the here studied solutions closely mirror, has for example been important for studying dualities in string theory Fateev et al. (2000); Alvarez et al. (1994), studying connections to little string theory Aharony et al. (1998); Maldacena and Strominger (1997), evaporation Callan et al. (1992); Russo et al. (1992), the Page curve Gautason et al. (2020), and holographic complexity Schneiderbauer and others (2020); Schneiderbauer et al. (2020). It would be interesting to further explore our solutions along those lines. More along the direction of studying Myers-Perry black holes, it can be insightful to use our results to, e.g., study black holes with different topologies and instabilities such as the Gregory-Laflamme type Gregory and Laflamme (1993). In particular a further study of the here presented black holes possessing multi angular momentum-like charges could reveal insights into dynamical features such as (in-)stability.
@@ Acknowledgements.
We thank Vyshnav Mohan, Bo Sundborg, and Lárus Thorlacius for stimulating discussions in the context of this article. We want to especially thank Roberto Emparan for suggesting the current narrative of this work and insightful comments. We also thank Gerben Oling for valuable comments on the draft. The work of WS is supported by a Starting Grant 2023-03373 from the Swedish Research Council and the Olle Engkvists Stiftelse. PT is supported by a Nordita visiting PhD student fellowship, Primus grant PRIMUS/23/SCI/005 and GAUK 425425 from Charles University.
@@ Appendix A Alternative derivation for rotation and temperature
We assume the following near-horizon form: